OpenStudy (anonymous):

The sum of three integers is 182. The sum of the first and second integers exceeds the third by 28. The third integer is 25 less than the first. Find the 3 integers.

OpenStudy (anonymous):

Given: \[a+b+c=182\] \[a+b=c+28\] \[c-25=a\] So, substitute away and you will find your integers.

OpenStudy (anonymous):

easy let x y z represent the integers x + y + z = 182 x + y = z + 28 z = 25 - x

OpenStudy (anonymous):

this makes no sense to me... :/

OpenStudy (anonymous):

| -52 | | 157 | | 77 __|